# Beyond time-homogeneity for continuous-time multistate Markov models

### v2

@@ Abstract
Keywords: Disease progression; Hidden Markov model; Hierarchical Bayesian modeling; Longitudinal study; State space model
@@ 1 Introduction
From a survival analysis perspective, it is understood that the Kolmogorov forward equations have a product integral solution expression for the transition probability matrix (see Andersen et al. (2012) and Borgan (1997) for more detail). An extension of the Kaplan-Meier estimator (Kaplan & Meier, 1958), the Aalen-Johansen estimator (Aalen & Johansen, 1978) is an empirical product limit estimator for the transition probability matrix of a finite state, inhomogeneous Markov process. An appeal of this estimator is that it can serve as a nonparametric MLE for the true transition probability matrix, and under certain conditions, is a uniformly consistent estimator (Andersen et al., 2012). Extending beyond Markov processes in particular, much work has been done to explore asymptotic properties of the Aalen-Johansen estimator in non-Markov settings (e.g., Datta & Satten, 2001, 2002; Glidden, 2002; Gunnes et al., 2007) as well as more recent work on incorporating covariate information into the evaluation of the Aalen-Johansen estimator (Bladt & Furrer, 2023). Although the Aalen-Johansen estimator offers a nonparametric estimation approach for the transition probability matrix, it does not estimate an exact solution for the transition probability matrix. Whether using the product integral or Reimann integral representation, the challenge of writing the transition probability matrix in closed-form for a continuous-time, time-inhomogeneous multistate Markov process still exists (Titman & Sharples, 2010; Titman, 2011; Cook & Lawless, 2014; Meng et al., 2022). Furthermore, focusing specifically on HMMs, it is not clear how to approximate the transition probability matrix using the Aalen-Johansen estimator. The justification is as follows: it is known that the Kaplan-Meier estimator can serve as a nonparametric proxy for assessing model goodness-of-fit when implementing other approaches to fitting multistate Markov models (Bureau et al., 2003; Titman & Sharples, 2008, 2010). Consequently, in Soper et al. (2020), the Kaplan-Meier approach to evaluating model fit is attempted when an HMM is fit to data; however, it is noted that these estimators are not “true” Kaplan-Meier estimators because there exists no approach for calculating these estimates while accounting for misclassification errors for the observed state labels of the Markov process. This logic can then be extended in the case of the Aalen-Johansen estimator. This offers, however, an interesting area for future research to explore how to calculate the Aalen-Johansen estimator under the framework of a latent state Markov process.
@@
More recently, the methods of Titman (2011) have been extended to accommodate the HMM framework in the R package nhm (Titman, 2019), and for some model specifications first derivatives of the likelihood function can be computed analytically. Bayesian implementations are not yet available in the nhm package. In addition to investigating a Bayesian approach, our paper provides further clarity, intuition, and empirical studies to illustrate the consequences of a piecewise-homogeneous approximation when transition rates may be continuous functions of time. Hence, our paper goes beyond the scope of what is presented in Titman (2011) by facilitating broader awareness for implementing these numerical ODE solvers within the context of continuous-time, time-inhomogeneous Markov models, as well as applying them to more complex data applications that go beyond the purview of what is capable of nhm. Additionally, Siméon et al. (2021) points to the lack of these numerical integration techniques being applied in the statistics literature for correctly fitting multistate Markov models (HMMs or otherwise). Furthermore, Siméon et al. (2021) argues that piecewise-homogeneous approximations are slow and inaccurate relative to numerical methods.
@@
An alternative to the continuous-time, piecewise-homogeneous approximation is to consider a discretization of time, altogether. A detailed comparison of discrete-time versus continuous-time, time-homogeneous models can be found in Wan et al. (2016). They find that if inter-observation times are similar, a discrete-time model is as effective as a continuous-time model, as well as computationally more efficient. See also, Aalen et al. (1997) and Liu et al. (2015) for discretized approaches for fitting continuous-time Markov models. We present evidence, in Section 3.4.2, suggesting the reasonableness of discrete-time approximations for time-inhomogeneous HMMs – when regularly-spaced observation times are available at a high resolution – based on our analysis of real electrocorticography (ECoG) data recorded from a mice sleep study (the data are from Bojarskaite et al., 2020). Nonetheless, the importance of our investigation of continuous-time processes is built on the fact that many data sets do not exhibit observation times near enough a regular spacing that a discrete-time approximation would be effective.
@@
The remainder of the paper is structured as follows. In Section 2, the multistate Markov model and HMM frameworks are introduced, and mathematical intuition is provided for the potential biases of a piecewise-homogeneous approximation for continuously evolving time-inhomogeneous transition rates in the context of a simple linear regression illustration. Section 3 covers all details for the data analyses beginning with data descriptions (Section 3.1) and modeling approaches (Section 3.2). Then, Section 3.3 presents a simulation study of synthetic data generated from a continuous-time, time-inhomogeneous Markov process, to compare fitting an HMM with piecewise-homogeneous approximations versus the time-inhomogeneous numerical ODE solution. We consider comparisons of these approaches on real data sets in Section 3.4. The paper ends with concluding remarks in Section 4. The R code along with a workflow script file for reproducing all numerical results presented in this paper is available at https://ebkendall.github.io/research.html.
@@ 2 Multistate Markov model methods and inference
We first formalize the common notations and specifications for multistate Markov models. For any $t,h\geq 0$, let ${\bf\it P}(t,t+h)$ be the $m\times m$ transition probability matrix over the time period from $t$ to $t+h$, for a Markov process $X(t)\in\{1,\dots,m\}$, for some fixed, positive integer $m$. Further, denote the component in the $r^{\text{th}}$ row and $s^{\text{th}}$ column of ${\bf\it P}(t,t+h)$ as
$P_{r,s}(t,t+h)=P\{X(t+h)=s\mid X(t)=r\},$
which represents the probability of transitioning from state $r$ to state $s$ over the period of time from $t$ to $t+h$. The mathematical expression for ${\bf\it P}(t,t+h)$ is determined by whether time is measured as discrete or continuous. While the discrete-time approach leads to a simple expression (i.e., multiplying the one-time-point-ahead transition probability matrix $h$ times, where $t$ and $h$ take values from a discrete grid of points such as the nonnegative integers), it is not adequate for modeling data that are observed irregularly in time. Thus, it is the necessary complexities of the continuous-time approach that we focus on in this paper.
@@
This closed-form solution makes it uncomplicated to derive the joint probability mass function for a collection of $n$ independent state sequences ${\bf\it s_{i}}:=\{s_{i,1},\dots,s_{i,n_{i}}\}$, of length $n_{i}$, observed at time points $t_{i,1},\dots,t_{i,n_{i}}$, for $i\in\{1,\dots,n\}$:
@@ 2.1 Simple illustration of piecewise-constant approximation bias
For fixed $d>0$, so long as sufficient variation is observed in the $g_{d}(t_{1}),\dots g_{d}(t_{n})$ (i.e., there is an unbounded and heavy-tailed distribution of available time instances $t_{1},\dots,t_{n}$, for large $n$), the bias $\beta_{1}-b_{1,n}$ for the slope parameter will vanish for large $n$, as stated in Lemma 1. The bias $\beta_{0}-b_{0,n}$ for the baseline parameter, however, need not vanish for large $n$, as illustrated by Lemma 2. Figure 1 presents a graphical representation of the these two lemmas.
@@ 2.2 Hidden Markov models
In order to evince the broadness of scope in which our analyses and conclusions apply, we further extend the continuous-time Markov model framework with the more pragmatic assumption that the underlying state sequences ${\bf\it s_{i}}=\{s_{i,1},\dots,s_{i,n_{i}}\}$, for $i\in\{1,\dots,n\}$, are not fully observable. With the added complication of latent or observed-with-error state sequences, the continuous-time, time-inhomogeneous multistate HMM framework applies. The same issues and potential rate biases – as discussed for (non-latent) Markov models – apply for approximating a time-inhomogeneous HMM by a piecewise-homogeneous HMM, but the likelihood function must be modified to account for uncertainty in the latent state sequences using some emitted response variable observations ${\bf\it y_{i}}=\{y_{i,1},\dots,y_{i,n_{i}}\}$, at time instances $t_{i,1},\dots,t_{i,n_{i}}$, for $i\in\{1,\dots,n\}$.
The standard HMM formulation assumes that the random variables $Y_{i,1},\dots,Y_{i,n_{i}}$ are conditionally independent given $s_{i,1},\dots,s_{i,n_{i}}$, and that $Y_{i,k}$ only depends on $s_{i,k}$, for every $k\in\{1,\dots,n_{i}\}$, for every $i\in\{1,\dots,n\}$. Then the density for ${\bf\it y_{i}}$ can be derived by integrating over all possible state sequences:
$f_{y_{i}}({\bf\it y_{i}})$ $=\sum_{s_{i,1}=1}^{m}\cdots\sum_{s_{i,n_{i}}=1}^{m}f_{y_{i}\mid s_{i}}({\bf\it y_{i}}\mid{\bf\it s_{i}})\cdot f_{s_{i}}({\bf\it s_{i}})=\sum_{s_{i,1}=1}^{m}\cdots\sum_{s_{i,n_{i}}=1}^{m}f_{s_{i}}({\bf\it s_{i}})\cdot\prod_{k=1}^{n_{i}}f_{y_{i,k}\mid s_{i,k}}(y_{i,k}\mid s_{i,k})$
@@ 3.1 Data descriptions
Three data sets are studied in Sections 3.3 and 3.4: real cardiac allograft vasculopathy (CAV) data (Section 3.1.1), synthetic CAV data (Section 3.1.2), and real ECoG mice sleep study data (Section 3.1.3). Each data application serves a unique purpose in illustrating the importance of using numerical ODE solvers for evaluating the transition probability matrix versus relying upon the piecewise time-homogeneous approximation, in the context of continuous-time, time-inhomogeneous Markov models. The focus is on comparing the resulting inferences from using the two aforementioned modeling approaches, and specifically how the estimation of the transition rate parameters differ.
Although common in the statistical literature to use the piecewise time-homogeneous assumption, if, for the real data analyses, the resulting parameter estimates are different from that of the numerical ODE approach, then the decision to use one method over the other is not to be overlooked. This is the motivation for the real data analyses because it is unknown which approach “truly” describes the process being modeled; therefore, if the inferential conclusions are different, practitioners must consciously decide if the piecewise-homogeneity assumption accurately represents the process, otherwise, the numerical ODE solver eliminates the need for that assumption. More broadly, the choice to use the CAV data is both because it is well-studied in the multistate Markov model literature, as well as it offers data with unequal inter-observation times within subjects; alternatively, the ECoG mice sleep study provides an example where the inter-observation times within and between subjects is the same. We generate synthetic CAV data in the context of the simulation study to illustrate how fitting the rate parameters of a continuous-time, time-inhomogeneous Markov process, based on a piecewise-homogeneous approximation – with rate matrix ${\bf\it Q}(g_{d}(t))$ – can lead to biased estimation of the baseline rates of ${\bf\it Q}(t)$.
@@ 3.1.2 Simulated data
The synthetic data is generated from a time-inhomogeneous Markov process based on HMM parameters fit to the well-studied CAV data set described in Section 3.1.1. In particular, the evolution of the Markov process rate matrix for this simulation is constructed to be inhomogeneous in continuous-time. Therefore, when the continuous-time, time-inhomogeneous HMM is fit to the synthetic data using both the piecewise-homogeneous and numerical ODE solution approaches, we are studying the model when it is both misspecified and correctly specified, respectively.
@@ 3.2 Competing methods
The four modeling approaches that will be compared are:
@@ 3.2.1 Remarks on computation
For the Bayesian approaches (A) and (C), the MCMC sampling is defined by a Metropolis-Hastings random walk. An adaptive proposal strategy is used during the burnin period to learn the proposal covariance for better mixing and quicker convergence. For approaches (B) and (D), maximum likelihood estimation is done via the BHHH (Titman, 2019) and BFGS (Jackson, 2011) algorithms, respectively. Lastly, both our custom numerical approach, (A), and the numerical approach in nhm, (B), implement the differential equation solvers from deSolve (Soetaert et al., 2010) which utilize the efficient and adaptive integration scheme developed in Petzold (1983). Note that when applying the nhm package, upper bounds on the parameter space are needed and successive observations with a time gap less than .001 are manually increased by .001 for numerical stability purposes.

### v3

@@ Abstract
Keywords: Hidden Markov model; Hierarchical Bayesian modeling; Longitudinal study; State space model; Aalen-Johansen estimator
@@ 1 Introduction
More recently, the methods of Titman (2011) have been extended to accommodate the HMM framework in the R package nhm (Titman, 2019), and for some model specifications first derivatives of the likelihood function can be computed analytically. Bayesian implementations are not yet available in the nhm package. In addition to investigating a Bayesian approach, our paper provides further clarity, intuition, and empirical studies to illustrate the consequences of a piecewise-homogeneous approximation when transition rates may be continuous functions of time. Hence, our paper goes beyond the scope of what is presented in Titman (2011) by facilitating broader awareness for implementing these numerical ODE solvers within the context of continuous-time, time-inhomogeneous Markov models, as well as applying them to more complex data applications that go beyond the purview of what is capable of nhm. Additionally, Siméon et al. (2021) points to the lack of these numerical-integration techniques being applied in the statistics literature for correctly fitting multistate Markov models (HMMs or otherwise). Furthermore, Siméon et al. (2021) argues that piecewise-homogeneous approximations are slow and inaccurate relative to numerical methods.
Another alternative to solving the time dependent Kolmogorov forward equations is to empirically estimate the transition probability matrix via the product-integral solution to the Kolmogorov forward equations (see Borgan (1997), Aalen et al. (2008), and Andersen et al. (2012) for more detail). An extension of the Kaplan-Meier estimator (Kaplan & Meier, 1958), the Aalen-Johansen estimator (Aalen & Johansen, 1978) is an empirical product limit estimator for the transition probability matrix of a finite state, inhomogeneous Markov process. Although the Aalen-Johansen estimator offers a nonparametric approach to estimating the transition probability matrix, the empirical calculation of the estimator assumes that all state transitions are observed and observation times correspond to state transition times (Borgan, 1997; Aalen et al., 2008). Hence, similar to the piecewise-homogeneous approach, the Aalen-Johansen estimator relies on making assumptions about the evolution of the state process which, in most applications, are likely not met and can thus result in estimation bias (as shown in Sections 3.3 and 3.4.1). That being stated, an appeal of this estimator is that it can serve as a nonparametric MLE for the true transition probability matrix, and under certain conditions, is a uniformly consistent estimator (Andersen et al., 2012). The Aalen-Johansen estimator is a function of the Nelson-Aalen estimator (Nelson, 1969, 1972), a nonparametric estimator of the cumulative transition intensity of the time-inhomogeneous Markov process. The Aalen-Johansen and Nelson-Aalen estimators are frequently used for assessing model goodness-of-fit when implementing other approaches to fitting multistate Markov models (Bureau et al., 2003; Titman & Sharples, 2008, 2010; Soper et al., 2020). Extending beyond Markov processes in particular, much work has been done to explore asymptotic properties of the Aalen-Johansen estimator in non-Markov settings (e.g., Datta & Satten, 2001, 2002; Glidden, 2002; Gunnes et al., 2007).
@@
The last alternative to solving the time dependent Kolmogorov forward equations we consider is a discretization of time, altogether. A detailed comparison of discrete-time versus continuous-time, time-homogeneous models can be found in Wan et al. (2016). They find that if inter-observation times are similar, a discrete-time model is as effective as a continuous-time model, as well as computationally more efficient. See also, Aalen et al. (1997) and Liu et al. (2015) for discretized approaches for fitting continuous-time Markov models. We present evidence, in Section 3.4.2, suggesting the reasonableness of discrete-time approximations for time-inhomogeneous HMMs – when regularly-spaced observation times are available at a high resolution – based on our analysis of real electrocorticography (ECoG) data recorded from a mice sleep study (the data are from Bojarskaite et al., 2020). Nonetheless, the importance of our investigation of continuous-time processes is built on the fact that many data sets do not exhibit observation times near enough a regular spacing that a discrete-time approximation would be effective.
@@
The remainder of the paper is structured as follows. In Section 2, the multistate Markov model and HMM frameworks are introduced, and mathematical intuition is provided for the potential biases of a piecewise-homogeneous approximation for continuously evolving time-inhomogeneous transition rates in the context of a simple linear regression illustration. Section 3 covers all details for the data analyses beginning with data descriptions (Section 3.1) and modeling approaches (Section 3.2). Then, Section 3.3 presents a simulation study of synthetic data generated from a continuous-time, time-inhomogeneous Markov process, to compare fitting an HMM with piecewise-homogeneous approximations, empirical approximations via the Aalen-Johansen and Nelson-Aalen estimators, as well as the time-inhomogeneous numerical ODE solution. We consider comparisons of these approaches on real data sets in Section 3.4. The paper ends with concluding remarks in Section 4. The R code along with a workflow script file for reproducing all numerical results presented in this paper is available at https://ebkendall.github.io/research.html.
@@ 2 Multistate Markov model methods and inference
We first formalize the common notations and specifications for multistate models. Let ${\bf\it s_{i}}=\{s_{i,1},\dots,s_{i,n_{i}}\}$, for $i\in\{1,\dots,n\}$ independent, represent an observed state sequence from a discrete, multistate process, $X(t)$, where for any time $t\geq 0$, $X(t)\in\{1,\dots,m\}$, for some fixed, positive integer $m$. The inferential goal with regard to these $n$ independent, multistate processes is to learn parameters characterizing the transitions between the $m$ states over time. If $t_{i,1}<t_{i,2}<\dots<t_{i,n_{i}}$, for $i\in\{1,\dots,n\}$, represents the observed time points for each ${\bf\it s_{i}}$, then the joint probability mass function is given by
$f_{s_{1},\dots,s_{n}}({\bf\it s_{1}},\dots,{\bf\it s_{n}})$ $=\prod_{i=1}^{n}P(X(t_{i,1})=s_{i,1})\cdot\prod_{k=2}^{n_{i}}P\quantity(X(t_{i,k})=s_{i,k}\mid\{X(t_{i,j})=s_{i,j}\}_{j=1}^{k-1}).$
For the remainder of this paper, however, we make the stochastic assumption that the evolution of $X(t)$ is defined by a Markov process. In particular, the Markov assumption simplifies the expression for $f_{s_{1},\dots,s_{n}}({\bf\it s_{1}},\dots,{\bf\it s_{n}})$ as follows. For any $t,h\geq 0$, let ${\bf\it P}(t,t+h)$ be the $m\times m$ transition probability matrix over the time period from $t$ to $t+h$, for $X(t)$, where the $r^{\text{th}}$ row and $s^{\text{th}}$ column of ${\bf\it P}(t,t+h)$, denoted $P_{r,s}(t,t+h)$, represents the probability of transitioning from state $r$ to state $s$ over the time period $t$ to $t+h$. The Markov assumption thus allows the joint probability mass function to be expressed as
$f_{s_{1},\dots,s_{n}}({\bf\it s_{1}},\dots,{\bf\it s_{n}})=\prod_{i=1}^{n}P(X(t_{i,1})=s_{i,1})\cdot P_{s_{i,1},s_{i,2}}(t_{i,1},t_{i,2})\cdots P_{s_{i,n_{i}-1},s_{i,n_{i}}}(t_{i,n_{i}-1},t_{i,n_{i}}),$ (1)
where
$P_{s_{i,k-1},s_{i,k}}(t_{i,k-1},t_{i,k})=P\{X(t_{i,k})=s_{i,k}\mid X(t_{i,k-1})=s_{i,k-1}\},$
for $k=2,3,\dots,n_{i}$. In what follows, we detail the various approaches for defining ${\bf\it P}(t,t+h)$ depending on further stochastic assumptions made about the evolution of $X(t)$.
The mathematical expression for ${\bf\it P}(t,t+h)$ is determined by whether time is measured as discrete or continuous. While the discrete-time approach leads to a simple expression (i.e., multiplying the one-time-point-ahead transition probability matrix $h$ times, where $t$ and $h$ take values from a discrete grid of points such as the nonnegative integers), it is not adequate for modeling data that are observed irregularly in time. Thus, it is the necessary complexities of the continuous-time approach that we focus on in this paper.
@@
This closed-form solution makes it uncomplicated to express (1) as
@@ 2.1 Simple illustration of piecewise-constant approximation bias
For fixed $d>0$, so long as sufficient variation is observed in the $g_{d}(t_{1}),\dots g_{d}(t_{n})$ (i.e., there is an unbounded and heavy-tailed distribution of available time instances $t_{1},\dots,t_{n}$, for large $n$), the bias $\beta_{1}-b_{1,n}$ for the slope parameter will vanish for large $n$, as stated in Lemma 1. The bias $\beta_{0}-b_{0,n}$ for the baseline parameter, however, need not vanish for large $n$, as illustrated by Lemma 2. Figure 1 presents a graphical representation of these two lemmas.
@@ 2.2 Hidden Markov models
In order to evince the broadness of scope in which our analyses and conclusions apply, we further extend the continuous-time Markov model framework with the alternative assumption that the state sequences ${\bf\it s_{i}}=\{s_{i,1},\dots,s_{i,n_{i}}\}$, for $i\in\{1,\dots,n\}$, are unobserved. The state sequences ${\bf\it s_{i}}$ are still assumed to be defined by a continuous-time, time-inhomogeneous Markov process; however, now let ${\bf\it y_{i}}=\{y_{i,1},\dots,y_{i,n_{i}}\}$ represent observed response data, where the standard HMM formulation assumes that their random variable analogues $Y_{i,1},\dots,Y_{i,n_{i}}$ are conditionally independent given the unobserved $s_{i,1},\dots,s_{i,n_{i}}$, and that $Y_{i,k}$ only depends on $s_{i,k}$, for every $k\in\{1,\dots,n_{i}\}$, for every $i\in\{1,\dots,n\}$. The same issues and potential rate biases – as discussed for (non-latent) Markov models – apply for approximating a time-inhomogeneous HMM by a piecewise-homogeneous HMM, but the likelihood function must be modified to account for uncertainty in the latent state sequences and the observed response data.
Given the standard HMM formulation, the density for ${\bf\it y_{i}}$ can be derived by integrating over all possible unobserved state sequences:
$f_{y_{i}}({\bf\it y_{i}})$ $=\sum_{s_{i,1}=1}^{m}\cdots\sum_{s_{i,n_{i}}=1}^{m}f_{y_{i}\mid s_{i}}({\bf\it y_{i}}\mid{\bf\it s_{i}})\cdot f_{s_{i}}({\bf\it s_{i}})=\sum_{s_{i,1}=1}^{m}\cdots\sum_{s_{i,n_{i}}=1}^{m}f_{s_{i}}({\bf\it s_{i}})\cdot\prod_{k=1}^{n_{i}}f_{y_{i,k}\mid s_{i,k}}(y_{i,k}\mid s_{i,k}),$
@@ 3.1 Data descriptions
Three data sets are studied in Sections 3.3 and 3.4: real cardiac allograft vasculopathy (CAV) data (Section 3.1.1), synthetic CAV data (Section 3.1.2), and real ECoG mice sleep study data (Section 3.1.3). Each data application serves a unique purpose in illustrating the practicality of using numerical ODE solvers for evaluating the transition probability matrix versus relying upon the piecewise time-homogeneous approximation or the Aalen-Johansen estimator, in the context of continuous-time, time-inhomogeneous Markov models. The focus is on comparing the resulting inferences from using the aforementioned modeling approaches, and specifically how the estimation of the transition rate parameters differ.
Although common in the statistical literature to use the piecewise time-homogeneous assumption, if, for the real data analyses, the resulting parameter estimates are different from that of the numerical ODE approach, then the decision to use one method over the other is not to be overlooked. This is the motivation for the real data analyses because it is unknown which approach “truly” describes the process being modeled; therefore, if the inferential conclusions are different, practitioners must consciously decide if the piecewise-homogeneity assumption accurately represents the process, otherwise, the numerical ODE solver eliminates the need for that assumption. The same applies for implementing the Aalen-Johansen estimator and the assumption of observing all state transition times. More broadly, the choice to use the CAV data is both because it is well-studied in the multistate Markov model literature, as well as it offers data with unequal inter-observation times within subjects; alternatively, the ECoG mice sleep study provides an example where the inter-observation times within and between subjects is the same. We generate synthetic CAV data in the context of the simulation study to illustrate how fitting the rate parameters of a continuous-time, time-inhomogeneous Markov process, based on a piecewise-homogeneous approximation – with rate matrix ${\bf\it Q}(g_{d}(t))$ – can lead to biased estimation of the baseline rates of ${\bf\it Q}(t)$.
@@ 3.1.2 Simulated data
The synthetic data is generated from a time-inhomogeneous Markov process based on HMM parameters fit to the well-studied CAV data set described in Section 3.1.1. In particular, the evolution of the Markov process rate matrix for this simulation is constructed to be inhomogeneous in continuous-time. Additionally, observations are made at irregular time instances, similar to the real CAV data, thus not guaranteeing all state transitions are observed, nor do the observation times necessarily correspond to transition times. Therefore, when the continuous-time, time-inhomogeneous HMM is fit to the synthetic data using either the piecewise-homogeneous approximation or the Aalen-Johansen estimator, compared to the numerical ODE solution approach, we are studying the model when it is both misspecified and correctly specified, respectively.
@@ 3.2 Competing methods
The five modeling approaches that will be compared are:
@@
(E) :
The empirical approximation to the transition rate and transition probability matrices using the Nelson-Aalen and Aalen-Johansen estimators, respectively (computed using the AalenJohansen package (Bladt & Furrer, 2023) for R).
@@ 3.2.1 Remarks on computation
For the Bayesian approaches (A) and (C), the MCMC sampling is defined by a Metropolis-Hastings random walk. An adaptive proposal strategy is used during the burnin period to learn the proposal covariance for better mixing and quicker convergence. For approaches (B) and (D), maximum likelihood estimation is done via the BHHH (Titman, 2019) and BFGS (Jackson, 2011) algorithms, respectively. Both our custom numerical approach, (A), and the numerical approach in nhm, (B), implement the differential equation solvers from deSolve (Soetaert et al., 2010) which utilize the efficient and adaptive integration scheme developed in Petzold (1983). Note that when applying the nhm package, upper bounds on the parameter space are needed and successive observations with a time gap less than .001 are manually increased by .001 for numerical stability purposes. Lastly, in all data applications in Sections 3.3 and 3.4, we assume a parametric form for the transition rate matrix. Note that the Aalen-Johansen and Nelson-Aalen estimators are not designed to be applied in the modeling setups of Sections 3.3 and 3.4 because the inferential focus for those analyses is to fit a given parametric form of the transition rates. Nonetheless, although the estimators for approach (E) are nonparametric, Section 5.2 provides the details of a method we constructed for associating the transition rate parameters with the Nelson-Aalen estimator over all time points (a verification of the efficacy of this approach is provided in the supplemental materials). It is important to highlight that the AalenJohansen package offers the prodint function which numerically computes the product-integral representation of the transition probability matrix, analogous to deSolve with Riemann integration. Thus, a possible alternative to approach (A) would be to use prodint instead of deSolve; however, a more thorough application of approach (A) with prodint is not conducted because it would have been computationally infeasible given the MCMC structure and the size of the data used (for more information, see the supplemental materials). That being said, based on preliminary investigations, approach (A) with either prodint or deSolve should lead to similar results.